# LLM Evaluation Based on Aerospace Manufacturing Expertise: Automated Generation and Multi-Model Question Answering

### v1

@@ II-A Evolution of LLM Evaluation Frameworks
The assessment paradigm for LLMs has evolved from single-task metrics to multidimensional frameworks. Foundational works like BLEU score [6] established task-specific performance measurement. Cross-domain benchmarks such as MMLU [4] and HellaSwag [5] introduced systematic evaluation of multitask understanding through carefully curated question sets. Recent advancements like Scylla [12] propose dynamic adversarial testing frameworks, addressing limitations of static evaluation paradigms.
@@ II-B Domain Adaptation Challenges
  • Vertical benchmarking: Medical QA systems employ licensure exam formats (e.g., MedQA-CMT [7]) to simulate real-world diagnostic reasoning
  • Interpretability integration: Legal evaluations combine performance metrics with SHAP explanation frameworks [13]
@@ II-C AI Integration in Aerospace Manufacturing
    Operational layer: Deep reinforcement learning enables PHM-guided predictive maintenance [9]
@@
    Quality assurance: Multi-view vision systems achieve 99.2

### v2

@@ II-A Evolution of LLM Evaluation Frameworks
The assessment paradigm for LLMs has evolved from single-task metrics to multidimensional frameworks. Foundational works like BLEU score [6] established task-specific performance measurement. Cross-domain benchmarks such as MMLU [4] and HellaSwag [5] introduced systematic evaluation of multitask understanding through carefully curated question sets. Recent advancements like Scylla [10] propose a dynamic evaluation framework that quantitatively measures LLMs’ generalization capabilities across varying task complexities. Through assessment of both in-distribution and out-of-distribution performance across 20 tasks at 5 complexity levels, Scylla reveals a non-monotonic relationship between task complexity and generalization ability, identifying critical thresholds where models transition from genuine reasoning to memorization-based responses. This framework provides deeper insights into how model scale affects generalization capabilities across different complexity regimes.
@@ II-B Domain Adaptation Challenges
  • Vertical benchmarking: Medical QA systems employ multilingual evaluation frameworks (e.g., MedExpQA [11]) to assess LLMs’ medical knowledge comprehension and reasoning capabilities across different languages
  • Interpretability integration: Legal evaluations combine performance metrics with SHAP explanation frameworks [shap2022]
@@ II-C AI Integration in Aerospace Manufacturing
    Operational layer: Deep reinforcement learning enables probabilistic RUL-based predictive maintenance, reducing maintenance costs by 29.3% while preventing 95.6% of unscheduled maintenance through Monte Carlo dropout-based CNN prognostics [7]
@@
    Quality assurance: Multi-view vision systems for aircraft inspection demonstrate promising potential for automating defect detection in composites [8]